# Dynamic and Distributed Routing in IoT Networks based on Multi-Objective Q-Learning

### v2

@@ Abstract
IoT networks often face conflicting routing goals such as maximizing packet delivery, minimizing delay, and conserving limited battery energy. These priorities can also change dynamically: for example, an emergency alert requires high reliability, while routine monitoring prioritizes energy efficiency to prolong network lifetime. Existing works, including many deep reinforcement learning approaches, are typically centralized and assume static objectives, making them slow to adapt when preferences shift. We propose a dynamic and fully distributed multi-objective Q-learning routing algorithm that learns multiple per-preference Q-tables in parallel and introduces a novel greedy interpolation policy to act near-optimally for unseen preferences without retraining or central coordination. A theoretical analysis further shows that the optimal value function is Lipschitz-continuous in the preference parameter, ensuring that the proposed greedy interpolation policy yields provably near-optimal behavior. Simulations show that our approach adapts in real time to shifting priorities and achieves up to 80–90% lower energy consumption and more than 2–5× higher cumulative rewards and packet delivery compared to six baseline protocols. These results demonstrate significant gains in adaptability, delivery, and efficiency for dynamic IoT environments.
@@ I Introduction
Designing efficient routing protocols for IoT requires balancing multiple, often contradictory objectives. For instance, the most energy-efficient path may traverse unreliable links and lead to high packet loss, while the most reliable or lowest-latency path may drain batteries quickly. Moreover, the relative importance of these objectives is not fixed: as an example, during time-critical events, latency and reliability are prioritized, whereas under normal conditions, energy conservation becomes more important to extend network lifetime. Future IoT systems must therefore support dynamic preference adaptation, where routing policies can adjust to shifting priorities in real time [liu2017performance, Vaishnav2024].
Classical approaches to multiobjective optimization often rely on weighted-sum formulations, where objectives are scaled by fixed weights and combined into a single cost function. While this provides a Pareto front of trade-offs, it assumes that preferences remain static [miettinen1999nonlinear]. In real-world IoT systems, however, priorities may change unpredictably during operation, making static weighting schemes insufficient.
Traditional routing solutions therefore fall short in dynamic IoT settings. Algorithms such as Dijkstra-based methods are computationally expensive for resource-limited devices. Heuristic schemes (e.g., fuzzy logic, clustering, or trust-based protocols) cannot adapt efficiently to changing objectives. Reinforcement learning (RL) approaches [kim2025deep, sharma2024optrisql, wang2024emor, 10419073, 10012574, 10608117, chakraborty2023intelligent, prabhu2023multiple, 9819969, 9488736, 9685191, 8935210] have emerged as promising alternatives, as they can learn routing decisions from interaction with the environment. However, most existing RL-based works (including deep RL methods) either (i) assume static objectives and retrain policies when preferences change, or (ii) rely on centralized controllers (e.g., SDN-based designs), which face problems of scalability, privacy, and single points of failure [frikha2021reinforcement]. This highlights the need to develop novel RL-based approaches for dynamic large-scale IoT deployments.
@@ I-A Motivation
Real-world IoT deployments require routing protocols that are dynamic, distributed, and resource-aware. Specifically:
  • Dynamic adaptation: Objectives such as packet delivery and energy efficiency change in importance depending on the network state and application needs. A protocol must adapt quickly without retraining from scratch.
  • Efficient knowledge reuse: Since retraining is costly and slow, the system must leverage past learning to adapt instantly to new preference settings.
  • Distributed operation: Each IoT node has limited energy and knowledge of the network. Protocols must work without global controllers or complete topology knowledge, using only local neighbor information.
These requirements motivate the design of a routing scheme that can simultaneously prepare for multiple objective trade-offs, act near-optimally for unseen dynamic preferences, and operate in a fully distributed manner.
@@ I-B Contributions and Novelty
This paper proposes a Dynamic Preference Q-Learning (DPQ-Learning) routing protocol that addresses the above challenges.
The key contributions of this work are summarized as follows:
@@
    Distributed Multi-Objective Q-Learning Framework: A fully distributed tabular Q-learning scheme that learns multiple per-preference Q-tables in parallel from a single behavior stream, enabling efficient reuse of experience under IoT constraints.
@@
    A Novel Greedy Interpolation Policy for Dynamic Preferences: A lightweight interpolation mechanism that allows seamless adaptation to sudden changes in preference weights without retraining or centralized control.
@@
    Theoretical Guarantee of Near-Optimality: A formal analysis establishing that the interpolated policy remains provably near-optimal across the entire range of preferences, providing theoretical support for the algorithm’s real-time adaptability and stability.
@@ II Related Works
[kim2025deep] DRL + Intelligent neighbor node selection - - - -
[sharma2024optrisql, chakraborty2023intelligent, farag2021congestion] Tabular QL Scalarised linear reward - - -
[wang2024emor] RL with Actor-Critic architecture RL-learned weights (energy vs. progress) Likely - - -
@@
[prabhu2023multiple] Multi-policy Deep-RL Ratio-based rewards - - - -
@@
[kaur2021energy] Deep-RL (DRL) Linear scalarization (LS) function - - -
@@
RL, among other data-driven ML approaches, has attracted special attention since it provides us with the facility to learn optimal policies while interacting with the environment. Thus, many RL-based routing protocols have been proposed [sharma2024optrisql, chakraborty2023intelligent, farag2021congestion, wang2024emor, prabhu2023multiple, kaur2021energy, kim2025deep, 10419073, 10608117, 8935210, abadi2022rlbeep, bouzid2020efficient, guo2019optimizing, 9819969, 9634122, 9488736, cong2021deep, ergun2022reinforcement, 9685191, 10012574, phan2022deepair]. A direct utilization of RL for IoT is to rely upon a central controller for route planning and control. This has been made possible by the recent development of software-defined networking (SDN). In SDN-enabled IoT networks, a central controller has complete knowledge of the network topology and takes most of the control decisions in the network. The SDN-based centralized protocols [8935210, 10419073, 9634122, phan2022deepair, 9488736, cong2021deep, 10012574, 9819969] face several challenges. Since the SDN controller plays a pivotal role in such protocols, it is an attractive target for malicious users to carry out DDoS attacks or other security attacks [javanmardi2023sdn]. Such routing protocols are also vulnerable to single-point failures. Another critical issue in IoT is maintaining individual IoT nodes’ data privacy. Sharing all the data with a central controller can be both cost-inefficient and also pose a risk to data privacy. Moreover, as the size of IoT networks is increasing rapidly, scalability is a major issue with these centralized protocols [cong2021deep].
@@
With a more intelligent organization, it is feasible and beneficial to employ distributed versions of RL algorithms for routing. As an example, in tabular Q-learning, SDN-based and other centralized approaches rely on a centralized controller to keep the knowledge of the entire network, store the entire Q-table, and do the whole route planning as the RL agent. However, the Q-table can be split into smaller parts, with each node having its local small Q-tables with information about and interaction with their immediate neighbors. In this way, each node acts as an RL agent capable of making its own decisions. Some existing research works utilize this possibility and thus propose distributed RL-based routing protocols [guo2019optimizing, bouzid2020efficient]. Abadi et al. [abadi2022rlbeep] utilize a sleep scheduling mechanism for enhancing the energy efficiency of their RL-based routing protocol. Such sleep scheduling is difficult to achieve in a fully distributed setup. Thus, they utilize a semi-distributed approach. They divide the network into clusters, with each cluster head acting as the local controller for the cluster. In this work, however, we showcase how a fully distributed RL-based scheme can be effectively applied while ensuring high energy efficiency and other vital objectives. Some other distributed RL-based routing protocols [sharma2024optrisql, chakraborty2023intelligent, farag2021congestion, kim2025deep] have also been proposed, but they lack in dynamically adapting to changing preferences, as described below (Comparisons are summarized in Table. I).
While the RL-based routing protocols discussed so far work well for the static objective preference on which the model is trained, they cannot deal with changing preferences and objectives. However, in IoT networks, the user and network preferences keep changing in runtime. Existing RL-based approaches cannot quickly adapt to changes in the environment, and it takes time to relearn new policies suitable for new preferences. For example, Natarajan et al. [natarajan2005dynamic] introduce an approach for distributed MORL with dynamic preferences. However, it does not approximate the entire Pareto front; rather, it learns from a limited number of random policies experienced in runtime. Further, it is limited to average reward RL algorithms (ARL) like R-learning and H-learning and not to discounted Q-learning. These approaches have some drawbacks that make them unsuitable for routing in IoT. It has been established in a previous work [mahadevan1994discount] that Q-learning consistently outperforms R-learning, even when both are evaluated with the same undiscounted performance metric. Additionally, R-learning is highly sensitive to the chosen exploration strategy. Interestingly, its performance notably worsens when using Boltzmann exploration. H-learning has the drawback that it consumes several times more CPU usage to converge than the model-free methods like Q-learning [tadepalli1994h], thus unsuitable for routing in IoT networks. To overcome all the above-mentioned challenges, we propose a novel multi-objective Q-Learning-based distributed routing scheme that can utilize the Pareto frontier learned to find efficient routing schemes swiftly adapting to the dynamically varying preferences in the network.
@@ IV-A Q-Learning
A key advantage of Q-learning, which we leverage in our approach, is that it is an off-policy algorithm [sutton1999reinforcement]. This means that the learning process is independent of the policy being followed during sample collection. In other words, Q-learning can learn the optimal Q-values regardless of the strategy used to explore the environment. As long as every state-action pair is explored sufficiently often, Q-learning is guaranteed to converge to the optimal Q-table. This property is particularly useful in multi-objective scenarios, since we can learn optimal Q-tables for many preferences simultaneously, irrespective of the actual routing policy being followed during the learning phase.
@@ IV-C Exploration Strategies
There are several approaches to efficiently balance the trade-off between exploration and exploitation. These methods typically focus on greater exploration in the early stages when little is known about the system. As learning progresses and the Q-tables approach the optimal values, more emphasis is placed on exploitation. In this paper, we examine two exploration strategies, though others can easily be integrated with our algorithm. The strategies we study are: (a) sequential exploration and exploitation, and (b) simultaneous exploration and exploitation.

### v4

@@ Abstract
IoT networks often face conflicting routing goals such as maximizing packet delivery, minimizing delay, and conserving limited battery energy. These priorities can also change dynamically: for example, an emergency alert requires high reliability, while routine monitoring prioritizes energy efficiency to prolong network lifetime. Existing works, including many deep reinforcement learning approaches, are typically centralized and assume static objectives, making them slow to adapt when preferences shift. We propose a dynamic and fully distributed multi-objective Q-learning routing algorithm that learns multiple per-preference Q-tables in parallel and introduces a novel greedy interpolation policy to act near-optimally for unseen preferences. The algorithm learns to optimize for energy efficiency, packet delivery ratio, and the composite reward, adapting to changing trade-offs between these metrics without retraining or centralized control. A theoretical analysis further shows that the optimal value function is Lipschitz-continuous in the preference parameter, ensuring that proposed greedy interpolation policy yields provably near-optimal behavior. Simulation results show that our approach adapts in real time to shifting priorities and achieves up to 80–90% lower energy consumption and up to 5 $\times$ higher cumulative rewards and packet delivery compared to six baseline protocols, under dynamic and distributed settings. Sensitivity analysis across varying preference window lengths confirms that the proposed DPQ framework consistently achieves higher composite reward than all baseline methods, demonstrating robustness to changes in operating conditions.
@@ I Introduction
Designing efficient routing protocols for IoT requires balancing multiple, often contradictory objectives. Moreover, the relative importance of these objectives is not fixed. For instance, sensor nodes may prioritize energy conservation during routine monitoring, selecting low-power routes that tolerate occasional packet loss. However, during a frost alert, the network may shift to prioritize a higher packet delivery ratio (PDR), possibly using higher-power links, redundant relays, or shorter hops, which can increase overall energy consumption. Future IoT systems must therefore support dynamic preference adaptation, where routing policies adjust to shifting priorities in real time without retraining [23, 34].
Classical approaches to multiobjective optimization often rely on weighted-sum formulations, where objectives are scaled by fixed weights and combined into a single cost function. While this provides a Pareto front of trade-offs, it assumes that preferences remain static [25]. In real-world IoT systems, however, priorities may change unpredictably during operation, making static weighting schemes insufficient.
Traditional routing solutions therefore fall short in dynamic IoT settings. Algorithms such as Dijkstra-based methods are computationally expensive for resource-limited devices. Heuristic schemes (e.g., fuzzy logic, clustering, or trust-based protocols) cannot adapt efficiently to changing objectives. Reinforcement learning (RL) approaches [19, 31, 36, 38, 14, 35, 5, 29, 37, 22, 9, 13] have emerged as promising alternatives, as they can learn routing decisions from interaction with the environment. However, most existing RL-based works (including deep RL methods) either (i) assume static objectives and retrain policies when preferences change, or (ii) rely on centralized controllers (e.g., SDN-based designs), which face problems of scalability, privacy, and single points of failure [11]. This highlights the need to develop novel RL-based approaches for dynamic large-scale IoT deployments.
@@ I-A Motivation
IoT routing must simultaneously balance multiple performance objectives such as packet delivery ratio (PDR) and energy consumption, while operating on resource-constrained devices. The relative priority of these objectives can differ across applications and may change over time based on network conditions or event-driven demands. A robust routing mechanism must therefore be flexible enough to accommodate varying objective trade-offs and computationally light enough to run in fully distributed IoT environments.
  • Dynamic adaptation to shifting priorities: IoT applications may require different performance goals at different times, for example, prioritizing PDR during critical events or prioritizing energy conservation during routine monitoring. Routing protocols must adapt to such changing preferences in real time.
  • Generalization across preferences without retraining: Current RL routing schemes typically learn policies tied to a fixed objective weighting and must be retrained when preferences change. This is slow and energy-intensive for IoT devices. A more efficient approach would allow reuse of past learning to accommodate new preference settings without re-learning policies from scratch.
  • Fully distributed and scalable operation for IoT: Many existing RL-based routing approaches rely on centralized control or global network information, which is impractical for IoT deployments where nodes operate autonomously with limited resources and only local neighbor knowledge.
These considerations motivate the development of a routing protocol that is dynamic, distributed, and able to adapt to changing objective preferences in real time. This work focuses on balancing three key IoT routing objectives: minimizing energy consumption, maximizing PDR, and capturing their context-dependent trade-offs through a composite reward.
@@ I-B Contributions and Novelty
This work introduces a fully distributed, dynamic multi-objective Q-learning routing protocol for IoT networks. The key contributions are:
@@
    Parallel per-preference Q-learning under IoT constraints: A lightweight tabular Q-learning scheme that learns a family of Q-tables for multiple objective preferences in parallel from a single behavior policy, enabling efficient reuse of experience.
@@
    A novel greedy interpolation policy (GIP) for unseen preferences: A distributed mechanism that allows each node to construct a routing policy for any unseen preference by interpolating between learned Q-values, enabling fast adaptation to dynamic objective trade-offs without retraining.
@@
    Theoretical guarantee of near-optimality: We establish analytical bounds showing that interpolated policies remains provably near-optimal across the entire range of preferences, with suboptimality controlled by the accuracy of learned Q-values and grid resolution.
[…]
@@ II Related Works
[19] DRL + Intelligent neighbor node selection - - - -
[31, 5, 10] Tabular QL Scalarised linear reward - - -
[36] RL with Actor-Critic architecture RL-learned weights (energy vs. progress) Likely - - -
@@
[29] Multi-policy Deep-RL Ratio-based rewards - - - -
@@
[18] Deep-RL (DRL) Linear scalarization (LS) function - - -
@@
RL, among other data-driven ML approaches, has attracted special attention since it provides us with the facility to learn optimal policies while interacting with the environment. Thus, many RL-based routing protocols have been proposed [31, 5, 10, 36, 29, 18, 19, 38, 35, 13, 1, 3, 12, 37, 4, 22, 6, 8, 9, 14, 28]. A direct utilization of RL for IoT is to rely upon a central controller for route planning and control. This has been made possible by the recent development of software-defined networking (SDN). In SDN-enabled IoT networks, a central controller has complete knowledge of the network topology and takes most of the control decisions in the network. The SDN-based centralized protocols [13, 38, 4, 28, 22, 6, 14, 37] face several challenges. Since the SDN controller plays a pivotal role in such protocols, it is an attractive target for malicious users to carry out DDoS attacks or other security attacks [17]. Such routing protocols are also vulnerable to single-point failures. Another critical issue in IoT is maintaining individual IoT nodes’ data privacy. Sharing all the data with a central controller can be both cost-inefficient and also pose a risk to data privacy. Moreover, as the size of IoT networks is increasing rapidly, scalability is a major issue with these centralized protocols [6].
@@
With a more intelligent organization, it is feasible and beneficial to employ distributed RL algorithms for routing. As an example, in tabular Q-learning, SDN-based and other centralized approaches rely on a centralized controller to keep the knowledge of the entire network, store the entire Q-table, and do route planning as the RL agent. However, the Q-table can be split into smaller parts, with each node having its local small Q-tables with information about and interaction with their immediate neighbors. In this way, each node acts as an RL agent capable of making its own decisions. Some existing research works utilize this possibility and thus propose distributed RL-based routing protocols [12, 3]. Abadi et al. [1] utilize a sleep scheduling mechanism for enhancing the energy efficiency of their RL-based routing protocol. Such sleep scheduling is difficult to achieve in a fully distributed setup. Thus, they utilize a semi-distributed approach. They divide the network into clusters, with each cluster head acting as the local controller for the cluster. In this work, however, we showcase how a fully distributed RL-based scheme can be effectively applied while ensuring high energy efficiency and other vital objectives. Some other distributed RL-based routing protocols [31, 5, 10, 19] have also been proposed, but they lack in dynamically adapting to changing preferences, as described below (Comparisons are summarized in Table I).
The Routing Protocol for Low-Power and Lossy Networks (RPL) is a protocol standardized by the Internet Engineering Task Force (IETF) for Low-Power and Lossy Networks, and it is widely used in many Internet of Things (IoT) deployments. In RPL, routing decisions are primarily driven by the objective function, which determines parent selection based on one or more routing metrics. As a result, a growing body of work has proposed reinforcement learning–enhanced RPL variants that aim to improve the adaptability of this decision process. Most RL-based RPL schemes formulate routing as a single-objective problem by combining multiple metrics—such as packet delivery ratio (PDR), delay, and energy—into a scalar reward using fixed or heuristically chosen weights [10, 15, 16]. Even when weights are adjusted online, for example through learning-automata-based tuning (e.g., DDSLA-/LALA-RPL [15, 16]) or context-aware objective functions [10], the learned policy ultimately converges to a single operating point for a given application profile. Consequently, adapting to changing application requirements typically requires re-tuning reward weights or re-running the learning process.
Our proposed scheme overcomes this limitation by supporting dynamic preference changes without retraining. While no existing work demonstrates a fully dynamic, distributed, multi-objective Q-learning routing scheme implemented strictly within RPL’s standard Objective Function and rank computation on resource-constrained nodes, several studies provide strong supporting evidence for the feasibility of such an approach. Prior work has shown that Q-learning can be effectively used for parent selection in RPL-based routing [10, 21, 30], and other studies demonstrate that learning-based decision logic can be embedded within RPL objective functions without modifying the protocol’s control plane [20]. Building on these insights, our proposed dynamic multi-objective Q-learning scheme can be mapped to RPL’s Objective Function and rank computation, enabling learned parent selection while preserving RPL’s existing control messages and DODAG structure. Moreover, prior implementations and evaluations in constrained IoT environments, such as those using Contiki and the Cooja simulator, suggest that such learning-based extensions to RPL are practically feasible [30].
More generally, existing RL-based routing approaches struggle to adapt quickly to changes in preference settings, as they require time to relearn policies suited to new objectives. For example, Natarajan et al. [27] introduce a distributed multi-objective reinforcement learning (MORL) approach with dynamic preferences. However, this method does not approximate the full Pareto front and instead learns from a limited number of random policies encountered during runtime. Furthermore, it is restricted to average-reward reinforcement learning algorithms such as R-learning and H-learning and does not extend to discounted Q-learning. These limitations are important in IoT settings, where learning efficiency and computational simplicity are critical. Prior work has shown that Q-learning consistently outperforms R-learning, even when both are evaluated using the same undiscounted performance metric [24]. In addition, R-learning is highly sensitive to the choice of exploration strategy, with performance degrading significantly under Boltzmann exploration, while H-learning incurs substantially higher computational overhead and slower convergence compared to model-free methods such as Q-learning [33]. These properties make average-reward methods less suitable for routing in resource-constrained IoT networks.
To overcome the above-mentioned challenges, we propose a novel multi-objective Q-learning–based distributed routing scheme that can exploit the learned Pareto frontier to adapt efficiently and swiftly to dynamically varying preferences in IoT networks. Our approach learns multiple Q-functions corresponding to different preference vectors and introduces a greedy interpolation policy that enables near-optimal routing decisions for previously unseen preferences without restarting learning or relying on centralized coordination. Importantly, we provide formal performance guarantees by proving Lipschitz continuity of the optimal value function with respect to preferences, which yields explicit bounds on the sub-optimality of interpolated policies. To the best of our knowledge, existing RL-based and other routing proposals do not support preference-continuous policies or offer theoretical guarantees under dynamic preference shifts.
@@ IV-A Q-Learning
A key advantage of Q-learning, which we leverage in our approach, is that it is an off-policy algorithm [32]. This means that the learning process is independent of the policy being followed during sample collection. In other words, Q-learning can learn the optimal Q-values regardless of the strategy used to explore the environment. As long as every state-action pair is explored sufficiently often, Q-learning is guaranteed to converge to the optimal Q-table. This property is particularly useful in multi-objective scenarios, since we can learn optimal Q-tables for many preferences simultaneously, irrespective of the actual routing policy being followed during the learning phase.
@@ IV-C Exploration Strategies
There are several approaches to efficiently balance the trade-off between exploration and exploitation. These methods typically focus on greater exploration in early stages when little is known about the system. As learning progresses and the Q-tables approach optimal values, more emphasis is placed on exploitation. In this paper, we examine two exploration strategies, though others can be integrated with our algorithm. The strategies we study are: (a) sequential exploration and exploitation, and (b) simultaneous exploration and exploitation.
@@
IV-D Objectives and Routing Metrics
The proposed approach is tailored to meet real-world IoT routing needs and evaluates routing decisions based on three key objectives:
  • Energy Consumption: Each transmission consumes energy depending on the hop distance. Minimizing cumulative energy usage is critical for IoT nodes with limited battery life.
  • Packet Delivery Ratio (PDR): This represents the reliability of a path and is influenced by link quality, node availability, and hop count. Maximizing PDR is essential in scenarios with critical data delivery needs.
  • Composite Reward: A weighted combination of the above metrics based on user-defined preference parameter $\beta\in[0,1]$, allowing the system to dynamically prioritize energy efficiency or reliability during runtime.
While the learning framework is general, the algorithm is explicitly implemented for multi-hop IoT routing under realistic constraints: distributed operation, lossy links, and battery-limited nodes. The preference-driven reward design and distributed learning scheme are tailored to unique challenges in IoT environments, distinguishing this work from generic routing strategies.
@@ V Distributed DPQ-Learning
Practical Adaptation to IoT Resource Constraints: The distributed DPQ-learning framework can be adapted to real-world IoT deployments by selecting the preference-grid size $|\mathcal{B}|$ and training budget according to per-node memory and compute limits. Since learning is localized, each node maintains only a local DPQ structure based on its one-hop neighborhood rather than storing Q-values for the full network. Let $|S_{\text{local}}|$ denote the number of local states represented at a node and $|A|$ be the number of candidate next-hop actions. Then, each preference point requires storing $|S_{\text{local}}|\cdot|A|$ Q-values, and the overall per-node storage is $|\mathcal{B}|\cdot|S_{\text{local}}|\cdot|A|$ floating-point values. With 64-bit floating-point storage, the memory footprint is approximately $8\,|\mathcal{B}|\,|S_{\text{local}}|\,|A|$ bytes per node, enabling a tunable accuracy–resource trade-off.
@@ VI-A Proof of Theorem 1
VI-B Limitations and Complexity Analysis
The DPQ-Learning framework provides strong adaptability but introduces several computational and practical trade-offs:
  • Memory complexity: $O(n\cdot|S|\cdot|A|)$
    The algorithm maintains one Q-table for each preference grid value, leading to memory usage that grows linearly with the number of grid points $n=|\mathcal{B}|$. This is manageable for moderate $n$ but may limit deployment on ultra-low-power IoT nodes.
  • Per-step update complexity: $O(n)$
    At each interaction step, all $n$ Q-tables are updated in parallel using the same transition sample. Although each update is lightweight, the total per-step cost increases linearly with $n$.
  • Action-selection overhead: $O(|A|)$
    The greedy interpolation policy requires evaluating two Q-values per action from neighboring grid points. Since interpolation is constant-time, the action-selection cost remains $O(|A|)$, similar to single-objective Q-learning.
  • Dependence on grid resolution
    Finer grids improve interpolation fidelity but incur higher memory and update costs, whereas coarser grids reduce resource usage at the expense of approximation error. This creates a tunable accuracy–complexity trade-off.
These considerations emphasize that the grid size and learning configuration should be tuned according to the computational constraints of the target IoT deployment. It should be noted that in reinforcement learning—particularly in multi-objective settings—the total runtime is inherently stochastic and depends on environment interactions, episode lengths, and convergence behavior. Thus, a closed-form time-complexity expression for the overall learning process is not provided. As already captured in the above discussion through memory requirements, per-step update costs, action-selection overhead, and grid-resolution trade-offs, computational performance is therefore best characterized via empirical runtime and sample efficiency rather than analytical bounds.